# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

### v3

@@ Abstract.
We prove a semi-global existence and convergence theorem for the $(3+1)$-dimensional vacuum Einstein equations with positive cosmological constant on spacetimes $\widetilde{M}\cong M\times\mathbb{R}$, where $M$ is a closed, connected, oriented three-manifold of negative Yamabe type. In constant mean curvature transported spatial coordinates, we show that solutions arising from a class of arbitrarily large initial data converge to a Riemannian metric of constant negative scalar curvature in infinite Newtonian-like ‘time’. As a consequence, the Einstein-$\Lambda$ flow generically fails to produce geometrization in the sense of Thurston. Our results affirm a conjecture of Ringström concerning the asymptotic in-distinguishability of spatial topology in the large data regime of Einstein-$\Lambda$ dynamics. A related result is established for positive Yamabe type under a technical condition.
@@ 1. Introduction
The six flat manifolds comprise by themselves the subset of zero Yamabe type. These admit metrics having vanishing scalar curvature (the flat ones) but no metrics having strictly positive scalar curvature. Here $\Gamma\subset SO(4)$ acts freely and properly discontinuously on $\mathbb{S}^{3}$. Also note that $M\#\mathbb{S}^{3}\approx M$ for any 3-manifold $M$. It is known that every prime $K(\pi,1)$ manifold is decomposable into a (possibly trivial but always finite) collection of (complete, finite volume) hyperbolic and graph manifold  [^2] A compact oriented $3$-manifold $M$ is a graph manifold if $\exists$ a finite disjoint collection of embedded 22-tori $\{T_{j}\}\subset M$ such that each connected component of $M-\cup_{j}T_{j}$ is the total space of circle bundle over surfaces components.
@@
1.1. Background Literature
The future stability problem for cosmological solutions to the Einstein equations has been extensively studied in the small-data regime, particularly when the spatial manifold MM lies in the negative Yamabe class and admits a hyperbolic metric.
@@ 1.2. Main Result
We consider the Einstein vacuum equations with positive cosmological constant $\Lambda>0$ on $(3+1)$-dimensional globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$, where $M$ is a closed, connected, oriented $3$-manifold of negative Yamabe type. Importantly, we do not assume that $M$ admits a negative Einstein metric; in particular, $M$ may contain graph manifold summands in its prime decomposition, cf. (2), for which no Einstein metric exists.
@@
which measures the failure of $g$ to be Einstein. In three spatial dimensions, by Mostow rigidity, $\mathfrak{T}[g]=0$ implies that $g$ is a hyperbolic metric, uniquely determined up to isometry. For generic MM of negative Yamabe type, $\mathfrak{T}[g]\neq 0$ for all $g$.
@@
Theorem 1.1 (Semi-global existence: $\Lambda>0$, $\sigma(M)\leq 0$).
@@
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on MM of pointwise constant negative scalar curvature, in the sense that
@@
Note that the allowed initial data size $\mathcal{I}^{0}$ is modulated by the condition $\mathcal{I}^{0}e^{-a/10}<1$ for initial time $T_{0}=a>0$ i.e., when $a$ is small, then the allowed data size is $O(1)$ but when $a\gg 1$, $\mathcal{I}^{0}$ can be very large.
@@
Remark 4.
This reproduces the elementary fact that in dimension $3$, there is an infinite-dimensional family of constant negative scalar curvature metrics that are non-conformal and non-isometric to the standard metrics on closed hyperbolic manifolds. Note that standard metrics are scalar curvature rigid in the conformal class i.e., conformal deformations of finite volume hyperbolic metrics preserving scalar curvature are necessarily isometries.
@@ 1.3. Heuristics on the genericity of the initial data
We comment on the notion of “genericity” of the initial data from a heuristic yet technically relevant viewpoint. The compact spatial manifold MM under consideration is of negative Yamabe type and, in particular, admits no Riemannian metric with nontrivial global Killing fields. Indeed, any compact 33–manifold that supports a continuous isometry group necessarily admits an effective SO(2)SO(2)–action. The classification theorem of Fischer–Moncrief [18] implies that no nontrivial connected sum of negative Yamabe type supports such an action. Among prime manifolds, only a distinguished subclass of Seifert fibered spaces—specifically those whose fundamental group has nontrivial infinite cyclic center—can admit an $SO(2)$ symmetry. Even in these exceptional geometries, such continuous symmetries arise only in a nongeneric subset of the moduli space of initial data. Thus, from the perspective of geometric constraints, the absence of continuous symmetries is the generic situation.
Although the topology of $M$ allows for considerable freedom in the selection of initial data, the data are not arbitrary from the viewpoint of their quantitative concentration properties. To analyze this, we introduce a heuristic amplitude–wavelength characterization of the pair $(\Sigma,\mathfrak{T})$, where $\Sigma$ denotes the trace–free part of the second fundamental form and $\mathfrak{T}$ denotes the obstruction tensor. We let
$(A_{\Sigma},L_{\Sigma}),\qquad(A_{\mathfrak{T}},L_{\mathfrak{T}})$
denote the respective amplitude–wavelength pairs of $\Sigma$ and $\mathfrak{T}$. Interpreting $(\Sigma,\mathfrak{T})$ as wave-type fields, we view $L_{\Sigma}$ and $L_{\mathfrak{T}}$ as characteristic length scales, while $A_{\Sigma}$ and $A_{\mathfrak{T}}$ denote dimensionally rescaled amplitudes.
The control of the initial data norm in Theorem 1.1 implies, at the highest order, the constraint
(7) $L_{\Sigma}^{-3}|A_{\Sigma}|^{2}\;+\;L_{\Sigma}^{-1}|A_{\Sigma}|^{2}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}^{3}\;\lesssim\;\big(\mathcal{I}^{0}\big)^{2},$
whereas for the lower-order norms one obtains
(8) $L_{\Sigma}^{-1}|A_{\Sigma}|^{2}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}\;+\;|A_{\Sigma}|^{2}L_{\Sigma}^{3}\;\lesssim\;e^{-2a}\big(\mathcal{I}^{0}\big)^{2}.$
Analogously, for the obstruction tensor $\mathfrak{T}$,
(9) $L_{\mathfrak{T}}^{-1}|A_{\mathfrak{T}}|^{2}\;+\;|A_{\mathfrak{T}}|^{2}L_{\mathfrak{T}}\;+\;|A_{\mathfrak{T}}|^{2}L_{\mathfrak{T}}^{3}\;\lesssim\;\big(\mathcal{I}^{0}\big)^{2},$
and by the relative smallness condition stated in Theorem 1.1, one may in particular have
$\mathcal{I}^{0}=O\!\left(e^{\frac{a}{10}}\right).$
Observe that the presence of the term $L_{\Sigma}^{-3}|A_{\Sigma}|^{2}$ in (7) allows for substantial concentration provided the lower-order bound (8) remains satisfied. For instance, if
$A_{\Sigma}=O\!\left(e^{-\alpha a}\right),\qquad L_{\Sigma}=O\!\left(e^{-\beta a}\right),\qquad\text{with}\quad 2\alpha+3\beta=\tfrac{9}{5},$
then for $a>0$, the inequalities (7)–(8) are compatible. Thus, for $\Sigma$, high-frequency small-amplitude concentration is allowed, whereas large amplitude on very small scales is prohibited. In sharp contrast, (9) demonstrates that $\mathfrak{T}$ is significantly more flexible: one may heuristically take
$|A_{\mathfrak{T}}|^{2}=O\!\left(e^{\tfrac{a}{10}}\right),\qquad L_{\mathfrak{T}}=O\!\left(e^{-\tfrac{a}{10}}\right),$
so that large amplitude concentrated on small scales is admissible for $\mathfrak{T}$ when $a\gg 1$.
To interpret this from the viewpoint of free gravitational radiation, let us decompose the Weyl curvature into its electric and magnetic components relative to the unit timelike vector $T$:
$E:=W(T,\cdot,T,\cdot),\qquad B:={}^{*}W(T,\cdot,T,\cdot).$
From the Gauss equation,
$E=\mathfrak{T}+O\big(|\Sigma|^{2}\big),$
and hence one may allow initial data with
$|A_{E}|^{2}=O\!\left(e^{\tfrac{a}{10}}\right),\qquad L_{E}=O\!\left(e^{-\tfrac{a}{10}}\right),$
which corresponds to high-frequency, large-amplitude electric Weyl curvature (short-pulse type) even in the positive cosmological constant regime. By contrast, via the Codazzi identity,
$B=O\big(\nabla\Sigma\big),$
the magnetic part inherits the amplitude–wavelength restrictions of $\Sigma$, and thus only high-frequency, small-amplitude configurations are allowed for $B$ at the initial slice. See Figure 1 for a schematic illustration.
@@
$|A_{B}|=O\!\left(e^{\tfrac{a}{20}}\right),\qquad L_{B}=O\!\left(e^{-\tfrac{a}{10}}\right),$
then gravitational collapse could ensue despite the rapid expansion induced by the positive cosmological constant. This suggests that the class of initial data considered here carries genuine physical significance, rather than merely constituting a nonempty technical scenario. A detailed semi-global existence analysis, including the potential formation of trapped surfaces in the cosmological regime, is currently in progress.
Figure 1. A rough heuristic depicting the characterization of the smooth initial data considered in the article. The horizontal axis denotes the support size on the initial slice $(M,g_{0})$ while the vertical axis does so for the amplitude. $a$ is the free parameter considered here that can modulate the necessary amplitude-wavelength as to how big the initial data can be, dependent on the $CMC$ time $T_{0}=a$ of the initial Cauchy slice. While the electric part of spacetime Weyl curvature can have any data of the three types including the data of short pulse type (type 2) shown above, the magnetic part of the same cannot be of the short pulse type. Type (3) is a typical magnetic field (of spacetime Weyl curvature) configuration that is allowed in this study. Here, the initial re-scaled geometry $(M,g_{0})$ is assumed to have $O(1)$ diameter.

### v4

# A large data result for vacuum Einstein’s equations

@@ Abstract.
[…]
@@ 1. Introduction
The six flat manifolds comprise by themselves the subset of zero Yamabe type. These admit metrics having vanishing scalar curvature (the flat ones) but no metrics having strictly positive scalar curvature. Here $\Gamma\subset SO(4)$ acts freely and properly discontinuously on $\mathbb{S}^{3}$. Also note that $M\#\mathbb{S}^{3}\approx M$ for any 3-manifold $M$. It is known that every prime $K(\pi,1)$ manifold is decomposable into a (possibly trivial but always finite) collection of (complete, finite volume) hyperbolic and graph manifold  [^2] A compact oriented 33-manifold $M$ is a graph manifold if $\exists$ a finite disjoint collection of embedded 22-tori $\{T_{j}\}\subset M$ such that each connected component of $M-\cup_{j}T_{j}$ is the total space of circle bundle over surfaces components.
@@
$1.$ Background Literature
The future stability problem for cosmological solutions to the Einstein equations has been extensively studied in the small-data regime, particularly when the spatial manifold $M$ lies in the negative Yamabe class and admits a hyperbolic metric.
@@ 1.2. Main Result
We consider the Einstein vacuum equations with positive cosmological constant $\Lambda>0$ on $(3+1)$-dimensional globally hyperbolic spacetimes $\widetilde{M}\cong M\times\mathbb{R}$, where $M$ is a closed, connected, oriented 33-manifold of negative Yamabe type. Importantly, we do not assume that $M$ admits a negative Einstein metric; in particular, $M$ may contain graph manifold summands in its prime decomposition, cf. (2), for which no Einstein metric exists.
@@
which measures the failure of $g$ to be Einstein. In three spatial dimensions, by Mostow rigidity, $\mathfrak{T}[g]=0$ implies that $g$ is a hyperbolic metric, uniquely determined up to isometry. For generic $M$ of negative Yamabe type, $\mathfrak{T}[g]\neq 0$ for all $g$.
@@
Theorem 1.1 (Global Well-posedness: $\Lambda>0$, $\sigma(M)\leq 0$).
@@
Moreover, the solution $(g(T),\Sigma(T))$ converges in the $C^{\infty}$ topology, as $T\to\infty$, to a limiting Riemannian metric $\widetilde{g}$ on $M$ of pointwise constant negative scalar curvature, in the sense that
@@
Note that the allowed initial data size $\mathcal{I}^{0}$ is modulated by the condition $\mathcal{I}^{0}e^{-a/10}<1$ for initial time $T_{0}=a>0$ i.e., when $a$ is small, then the allowed data size is $O(1)$ but when $a\gg 1$, $\mathcal{I}^{0}$ can be very large $O(e^{a/10})$.
@@ 1.3. Heuristics on the genericity of the initial data
We comment on the notion of “genericity” of the initial data from a heuristic yet technically relevant viewpoint. The compact spatial manifold $M$ under consideration is of negative Yamabe type and, in particular, admits no Riemannian metric with nontrivial global Killing fields. Indeed, any compact 33–manifold that supports a continuous isometry group necessarily admits an effective SO(2)SO(2)–action. The classification theorem of Fischer–Moncrief [18] implies that no nontrivial connected sum of negative Yamabe type supports such an action. Among prime manifolds, only a distinguished subclass of Seifert fibered spaces—specifically those whose fundamental group has nontrivial infinite cyclic center—can admit an $SO(2)$ symmetry. Even in these exceptional geometries, such continuous symmetries arise only in a nongeneric subset of the moduli space of initial data. Thus, from the perspective of geometric constraints, the absence of continuous symmetries is the generic situation.
Although the topology of $M$ leaves substantial freedom in the choice of initial data, the hypotheses of Theorem 1.1 still impose a definite restriction on how such data may concentrate. To describe this heuristically, we associate to the pair $(\Sigma,\mathfrak{T})$ amplitude–wavelength parameters
$(A_{\Sigma},L_{\Sigma}),\qquad(A_{\mathfrak{T}},L_{\mathfrak{T}}),$
where $\Sigma$ is the trace-free part of the second fundamental form and $\mathfrak{T}$ is the renormalized trace-free Ricci tensor appearing in Theorem $1.$ The quantities $L_{\Sigma}$ and $L_{\mathfrak{T}}$ should be interpreted as characteristic spatial scales, while $A_{\Sigma}$ and $A_{\mathfrak{T}}$ represent dimensionally normalized amplitudes. This discussion is purely heuristic and is intended only to indicate the concentration regimes compatible with the norm bounds of the theorem.
In three spatial dimensions, a field of amplitude AA concentrated at wavelength $L$ has squared $H^{k}$-size of order
$|A|^{2}\sum_{j=0}^{k}L^{3-2j}.$
Consequently, the control required in Theorem 1.1 yields, at the highest order for $\Sigma$,
(7) $L_{\Sigma}^{-3}|A_{\Sigma}|^{2}+L_{\Sigma}^{-1}|A_{\Sigma}|^{2}+L_{\Sigma}|A_{\Sigma}|^{2}+L_{\Sigma}^{3}|A_{\Sigma}|^{2}\lesssim(\mathcal{I}^{0})^{2},$
whereas the lower-order weighted bounds for $\Sigma$ give
(8) $L_{\Sigma}^{-1}|A_{\Sigma}|^{2}+L_{\Sigma}|A_{\Sigma}|^{2}+L_{\Sigma}^{3}|A_{\Sigma}|^{2}\lesssim e^{-2a}(\mathcal{I}^{0})^{2}.$
Similarly, for $\mathfrak{T}$ one obtains
(9) $L_{\mathfrak{T}}^{-1}|A_{\mathfrak{T}}|^{2}+L_{\mathfrak{T}}|A_{\mathfrak{T}}|^{2}+L_{\mathfrak{T}}^{3}|A_{\mathfrak{T}}|^{2}\lesssim(\mathcal{I}^{0})^{2}.$
Moreover, the smallness assumption in Theorem 1.1 allows $\mathcal{I}^{0}$ to grow with $a$, subject only to
$\mathcal{I}^{0}e^{-a/10}\ll 1.$
In particular, one may heuristically regard $\mathcal{I}^{0}$ as being as large as $e^{a/10}$, up to a fixed small factor.
The contrast between (7) and (8) is important. The term
$L_{\Sigma}^{-3}|A_{\Sigma}|^{2}$
shows that $\Sigma$ may carry nontrivial high-frequency concentration at the top order, but only under the severe lower-order restriction imposed by (8). Indeed, if one writes
$A_{\Sigma}\sim e^{-\alpha a},\qquad L_{\Sigma}\sim e^{-\beta a},$
and takes $\mathcal{I}^{0}\sim e^{a/10}$, then compatibility of (7)–(8) requires, at the level of exponents,
$3\beta-2\alpha\leq\frac{1}{5},\qquad\beta-2\alpha\leq-\frac{9}{5}.$
Thus $\Sigma$ may oscillate on short scales, but only with correspondingly small amplitude. For example,
$A_{\Sigma}=O(e^{-a}),\qquad L_{\Sigma}=O(e^{-a/10})$
is admissible. In this sense, the theorem allows high-frequency, small-amplitude concentration for $\Sigma$, but excludes a genuine short-pulse regime for $\Sigma$ itself.
By contrast, (9) is markedly less restrictive. At the heuristic threshold allowed by $\mathcal{I}^{0}\sim e^{a/10}$, one may take
$|A_{\mathfrak{T}}|^{2}=O(e^{a/10}),\qquad L_{\mathfrak{T}}=O(e^{-a/10}),$
for which the left-hand side of (9) is of order
$e^{a/5}+1+e^{-a/5}.$
This is compatible with the bound $(\mathcal{I}^{0})^{2}\sim e^{a/5}$. Heuristically, the tensor $\mathfrak{T}$ may therefore exhibit a genuine short-pulse-type concentration: large amplitude supported at very small spatial scale.
This has a natural interpretation in terms of the free gravitational field. Relative to the future-directed unit normal $T$, let
$E:=W(T,\cdot,T,\cdot),\qquad B:={}^{*}W(T,\cdot,T,\cdot)$
denote the electric and magnetic parts of the Weyl tensor. The trace-free Gauss–Codazzi relations show schematically that
$E=\mathfrak{T}+\mathrm{l.o.t.}(\Sigma),\qquad B=\operatorname{curl}\Sigma,$
where $\mathrm{l.o.t.}(\Sigma)$ denotes expressions of lower differential order in $\Sigma$ (depending on the precise normalization, these may include terms linear and quadratic in $\Sigma$). Since the lower-order norms of $\Sigma$ are strongly suppressed, the electric part may inherit the short-pulse concentration carried by $\mathfrak{T}$, while the magnetic part remains constrained by the smallness of $\Sigma$ and its first derivative. Thus the class of data constructed here permits high-frequency, large-amplitude electric Weyl curvature, but only high-frequency, small-amplitude magnetic Weyl curvature on the initial slice. See Figure 1 for a schematic illustration, and Section 4 for the rigorous construction.
@@
1.4. Heuristics for the construction of the initial data
Recall that in CMC gauge the re-scaled vacuum constraint equations reduce to
(10) $R(g)+\frac{n-1}{n}$ $=|\Sigma|_{g}^{2},$
(11) $\nabla^{j}\Sigma_{ij}$ $=0,$
where $\Sigma$ is the trace-free part of the second fundamental form. Thus $\Sigma$ is, by definition, trace-free, and (11) is the momentum constraint.
For the class of data considered in Theorem 1.1, one seeks
$\|\Sigma\|_{H^{2}}\lesssim e^{-a}\mathcal{I}^{0},\qquad\|\Sigma\|_{H^{3}}\lesssim\mathcal{I}^{0},\qquad\|\mathfrak{T}[g]\|_{H^{2}}\lesssim\mathcal{I}^{0},$
with
$\mathcal{I}^{0}e^{-a/10}<1.$
At the heuristic upper threshold allowed by the theorem, one may think of
$\mathcal{I}^{0}\sim e^{a/10},\qquad a\gg 1.$
The Hamiltonian constraint (10) then forces the renormalized scalar-curvature defect to be much smaller than the $H^{2}$-size of $\mathfrak{T}[g]$. Indeed, since $H^{2}$ is an algebra in dimension three,
$\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{2}}=\||\Sigma|_{g}^{2}\|_{H^{2}}\lesssim\|\Sigma\|_{H^{2}}^{2}\lesssim e^{-2a}(\mathcal{I}^{0})^{2},$
while at one higher derivative one only obtains
$\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{3}}=\||\Sigma|_{g}^{2}\|_{H^{3}}\lesssim\|\Sigma\|_{H^{2}}\|\Sigma\|_{H^{3}}+\|\Sigma\|_{H^{3}}^{2}\lesssim(\mathcal{I}^{0})^{2}.$
If $\mathcal{I}^{0}\sim e^{a/10}$, these bounds become
$\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{2}}=O(e^{-9a/5}),\qquad\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{3}}=O(e^{a/5}).$
Accordingly, the geometric core of the construction is the following. One must produce a metric $g$ on a closed three-manifold of negative Yamabe type such that
$\|\mathfrak{T}[g]\|_{H^{2}}=O(e^{a/10}),\qquad\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{2}}=O(e^{-9a/5}),\qquad\Bigl\|R(g)+\frac{n-1}{n}\Bigr\|_{H^{3}}=O(e^{a/5}),$
while at the same time arranging for a trace-free tensor $\Sigma$ satisfying
$\|\Sigma\|_{H^{2}}=O(e^{-9a/10}),\qquad\|\Sigma\|_{H^{3}}=O(e^{a/10}),$
and the momentum constraint.
The standard conformal method provides the natural framework for solving (10)–(11), but in the present setting the delicate point is not merely the solvability of the constraints. Rather, one must solve them in such a way that the conformal correction needed to enforce (10) remains sufficiently small at low order so that the large $H^{2}$-size of the renormalized trace-free Ricci tensor is not destroyed.
The construction therefore proceeds heuristically in three steps. First, one builds a seed metric $\widehat{g}$ for which the renormalized trace-free Ricci tensor is large in $H^{2}$, whereas the scalar-curvature defect $R(\widehat{g})+\frac{n-1}{n}$ is very small in $H^{2}$ and controlled in $H^{3}$. Second, one constructs a $\widehat{g}-$transverse-traceless tensor $\widehat{\Sigma}$ with the quantitative bounds required in Theorem 1.1; in particular, this resolves the momentum constraint at the seed level. Third, one uses the conformal method to solve the Hamiltonian constraint exactly, producing a nearby pair $(g,\Sigma)$ satisfying the full system (10)-(11). One then shows that the conformal correction is sufficiently small in the relevant norms, so that the large $H^{2}$-size of the renormalized trace-free Ricci tensor is retained.
This is the geometric mechanism underlying the initial-data construction. The following discussion here is only heuristic; the rigorous implementation is carried out in Section 4.
Let $M$ be a closed connected 33-manifold of negative Yamabe type. By definition, every conformal class on $M$ contains a smooth metric of constant negative scalar curvature. We therefore fix a smooth background metric $\gamma$ satisfying
$R(\gamma)=-\frac{n-1}{n}.$
Since throughout this discussion $n=3$, one may equivalently write $R(\gamma)=-\frac{2}{3}$; for notational clarity, however, we retain the general $n$-dimensional form in the formulas below.
We begin with a smooth background metric $\gamma$ satisfying
$R[\gamma]=-\frac{n-1}{n}.$
We then consider a perturbation
$g_{1}=\gamma+\epsilon h,\qquad 0<\epsilon\ll 1,$
where $h$ is chosen to be transverse-traceless with respect to $\gamma$,
$\mbox{div }_{\gamma}h=0,\qquad\mbox{tr}_{\gamma}h=0,$
and normalized by
$\|h\|_{L^{2}(M,\gamma)}=1.$
The essential point is that $h$ is taken to be highly oscillatory, with characteristic frequency $\mu\gg 1$. Concretely, one may choose $h$ from a high-frequency family of transverse-traceless tensors associated with any self-adjoint elliptic operator on the transverse-traceless bundle whose principal symbol is that of the rough Laplacian. Standard elliptic theory then yields the scale of estimates
$\|h\|_{H^{2}(M,\gamma)}\sim\mu,$
with constants depending only on the background geometry. If the background happens to be Einstein, one may take $h$ to be a high-frequency transverse-traceless eigentensor of the Lichnerowicz Laplacian; however, that special identification is not needed for the argument below.
We write
$\mathfrak{T}[g]:=\text{Ric}[g]-\frac{1}{n}R[g]\,g$
for the trace-free Ricci tensor. The linearization of $\mathfrak{T}$ at $\gamma$, restricted to transverse-traceless directions, is a second-order elliptic operator; we denote it schematically by $\mathcal{A}_{\gamma}$. Its principal part is the Laplace-type term, and therefore it carries the same high-frequency scaling as a second-order elliptic operator. Accordingly, one has the schematic expansion
$\mathfrak{T}[g_{1}]=\mathfrak{T}[\gamma]+\epsilon\,\mathcal{A}_{\gamma}h+\epsilon^{2}\mathcal{Q}_{\gamma}(h,\nabla h,\nabla^{2}h),$
where $\mathcal{Q}_{\gamma}$ is quadratic in $h$ and its derivatives. At the heuristic level relevant for the introduction, the ellipticity of $\mathcal{A}_{\gamma}$ implies
$\|\mathcal{A}_{\gamma}h\|_{H^{2}(M,\gamma)}\sim\mu^{2},$
whereas the quadratic remainder obeys the scale
$\|\mathcal{Q}_{\gamma}(h,\nabla h,\nabla^{2}h)\|_{H^{2}(M,\gamma)}\lesssim\mu^{3}.$
Thus
$\|\mathfrak{T}[g_{1}]\|_{H^{2}}\gtrsim\epsilon\mu^{2}-\epsilon^{2}\mu^{3}-O(1).$
This already exhibits the mechanism we exploit: the linear contribution to the trace-free Ricci tensor gains two powers of frequency, while the nonlinear error is smaller provided $\epsilon\mu$ remains sufficiently small.
The scalar curvature behaves differently. Since $h$ is transverse-traceless, its first variation is only zeroth order:
$DR_{\gamma}[h]=-\langle\text{Ric}^{\circ}[\gamma],h\rangle_{\gamma},$
where $\text{Ric}^{\circ}[\gamma]=\mathfrak{T}[\gamma]$. Consequently, the scalar curvature expansion has the schematic form
$R[g_{1}]=R[\gamma]+\epsilon\,\langle\text{Ric}^{\circ}[\gamma],h\rangle_{\gamma}+\epsilon^{2}\mathcal{R}_{\gamma}(h,\nabla h,\nabla^{2}h),$
with $\mathcal{R}_{\gamma}$ quadratic. Since the linear term involves no second derivatives of $h$, it is smaller by one power of frequency than the linear term in the trace-free Ricci tensor. Correspondingly,
$\Big\|R[g_{1}]+\frac{n-1}{n}\Big\|_{H^{2}}\lesssim\epsilon\mu+\epsilon^{2}\mu^{3}.$
The two expansions therefore, separate the relevant scales:
$\|\mathfrak{T}[g_{1}]\|_{H^{2}}\sim\epsilon\mu^{2},\qquad\Big\|R[g_{1}]+\frac{n-1}{n}\Big\|_{H^{2}}\sim\epsilon\mu+\epsilon^{2}\mu^{3},$
up to lower-order terms. In particular, we seek parameters for which
$\epsilon\mu^{2}\gg 1,\qquad\epsilon\mu\ll 1,\qquad\epsilon^{2}\mu^{3}\ll 1.$
For the later quantitative argument, it is convenient to enforce the stronger bounds
$\epsilon^{2}\mu^{3}<e^{-a/10},\qquad\epsilon\mu^{2}>e^{a/10},$
for some large parameter $a\gg 1$. Solving these inequalities yields
$e^{a/20}\epsilon^{-1/2}<\mu<\epsilon^{-2/3}e^{-a/30}.$
Equivalently,
$\epsilon^{1/2}e^{a/20}<\epsilon\mu<e^{-a/30}\epsilon^{1/3}.$
This interval is nonempty provided
$\epsilon^{1/2}e^{a/20}<e^{-a/30}\epsilon^{1/3},$
that is,
$\epsilon^{1/6}<e^{-a/12},\qquad\text{equivalently}\qquad\epsilon<e^{-a/2}.$
Thus, once $\epsilon$ is chosen sufficiently small relative to $a$, one may select a transverse-traceless tensor of frequency $\mu$ in the admissible range
(12) $e^{a/20}\epsilon^{-1/2}<\mu<\epsilon^{-2/3}e^{-a/30}.$
At the heuristic level, this produces a metric $g_{1}$ for which the trace-free Ricci tensor is large in $H^{2}$, while the scalar-curvature defect remains perturbative. This, for the appropriate choice of $\epsilon<e^{-a/2}$, yields the bounds
[…]
However, the task is not completed since we need to construct the physical initial data $(g_{0},k_{0})$ that verifies the constraint equations. However, note that the momentum constraint is relatively easy to address. On the other hand, the Hamiltonian constraint deals with the scalar curvature, which, in principle, can be modified almost independently of the trace-free Ricci curvature by means of a conformal transformation. We do this in the second step.
The second step is to solve the momentum and Hamiltonian constraint through a conformal deformation. Starting from the metric $g=g_{1}$ constructed above, together with a symmetric transverse-traceless tensor $\Sigma$ relative to $g$, we seek a positive conformal factor $\varphi$ and define
$g_{0}=\varphi^{\frac{4}{n-2}}g,\qquad\Sigma_{0}=\varphi^{-2}\Sigma.$
Here $(g_{0},\Sigma_{0})$ would be the physical initial data set that we desire. The conformal covariance of the momentum constraint ensures that $\Sigma_{0}$ remains transverse-traceless with respect to $g_{0}$:
$\mbox{div }_{g_{0}}\Sigma_{0}=0$
[…]
(13) $R[g_{0}]=\varphi^{-\frac{n+2}{n-2}}\left(-\frac{4(n-1)}{n-2}\Delta_{g}\varphi+R[g]\varphi\right).$
Hence, the Hamiltonian constraint
$R[g_{0}]+\frac{n-1}{n}=|\Sigma_{0}|_{g_{0}}^{2}$
reduces to the usual Lichnerowicz equation for $\varphi$. For the purposes of the present discussion, the important point is simply that the conformal factor is chosen so as to remove the scalar-curvature defect left over from the first step.
What must still be verified is that this scalar-curvature correction does not significantly alter the large trace-free Ricci component already present in $g$. The transformation law for the trace-free Ricci tensor is
$\mathfrak{T}[g_{0}]=\mathfrak{T}[g]-(n-2)\varphi^{-1}\left(\nabla^{2}\varphi-\frac{1}{n}\Delta_{g}\varphi\,g\right)+\frac{2(n-1)}{n-2}\varphi^{-2}\left(d\varphi\otimes d\varphi-\frac{1}{n}|\nabla\varphi|_{g}^{2}g\right).$
The detailed estimates established later show that the correction term on the right-hand side is negligible in $H^{2}$ in light of $H^{4}$ norm of $\varphi$ being close to 1:
$\left\|-(n-2)\varphi^{-1}\left(\nabla^{2}\varphi-\frac{1}{n}\Delta_{g}\varphi\,g\right)+\frac{2(n-1)}{n-2}\varphi^{-2}\left(d\varphi\otimes d\varphi-\frac{1}{n}|\nabla\varphi|_{g}^{2}g\right)\right\|_{H^{2}}=O(e^{-a}).$
Therefore
$\|\mathfrak{T}[g_{0}]-\mathfrak{T}[g]\|_{H^{2}}=O(e^{-a}),$
so the large $H^{2}$-size of the trace-free Ricci tensor survives the conformal correction. In particular,
$\|\mathfrak{T}[g_{0}]\|_{H^{2}}\sim\|\mathfrak{T}[g]\|_{H^{2}},$
up to an error which is negligible on the scale relevant to the construction. Therefore, we end up with the desired open set of initial data set $(g_{0},\Sigma_{0})$ that verifies
$R[g_{0,a}]+\frac{2}{3}$ $=|\Sigma_{0,a}|_{g_{0,a}}^{2},$
$\mbox{div }_{g_{0,a}}\Sigma_{0,a}$ $=0,$
$\mbox{tr}_{g_{0,a}}\Sigma_{0,a}$ $=0,$
and
[…]
In summary, the initial data are produced by combining two mechanisms with complementary effects. A high-frequency transverse-traceless perturbation creates a large trace-free Ricci component while keeping the scalar curvature nearly constant, and a subsequent conformal deformation restores the Hamiltonian constraint while modifying the trace-free Ricci tensor only perturbatively. Simultaneously one explicitly constructs the TTTT tensor $\Sigma$ verifying large $H^{3}$ norm while small $H^{2}$ norm (in appropriate scale). This is the geometric content behind the construction; the rigorous implementation is carried out in Section 4.
Figure 1. A rough heuristic depicting the characterization of the smooth initial data considered in the article in case the spatial slice $M$ admits a hyperbolic metric. This clearly depicts the difference between the current work and previous work in this framework. The red center is the isolated hyperbolic fixed point (Mostow rigidity). The internal yellow ball $D_{\delta}(\gamma,0)$ depicts the small data regime studied in [13, 34] where $d_{0}$ is the $H^{s}\times H^{s-1}$ Sobolev distance for $s>3/2+1$. In contrast current study can handle data lying within the ball $B_{a}((\gamma,0),\mathcal{I}^{0})$ where the ball $B_{a}$ is with respect to the distance as defined in (277) of the main theorem. This ball can be made arbitrarily large by increasing the parameter $a$.